OpenStudy (phebe):

SOME HELPP HERE: WILLL GIVVV MEDALL!! 3. Put these numbers in order from least to greatest: 1.2 x 103; 2.5 x 10-2; 2.8 x 10-3; 4 x 103

OpenStudy (phebe):

@Callisto

OpenStudy (phebe):

can u help me with this @Callisto

OpenStudy (callisto):

Are the numbers \[1.2\times 10^3; 2.5 \times 10^{-2}; 2.8 \times 10^{-3}; 4 \times 10^3\]?

OpenStudy (phebe):

yess

OpenStudy (phebe):

@Callisto

OpenStudy (callisto):

First, arrange then according to the power of 10, can you do it?

OpenStudy (phebe):

yes but um. do u know the answerr

OpenStudy (callisto):

Then, do it first. You'll get the answer soon.

OpenStudy (phebe):

ok but im behiond nd im tryin to hurry so nvm then k

OpenStudy (phebe):

@Imad_Sawal u think u can help me

OpenStudy (anonymous):

Yup .

OpenStudy (anonymous):

Ok , so we've got 1.2 x 10^3; 2.5 x 10^2; 2.8 x 10^3; 4 x 10^3

OpenStudy (anonymous):

4 numbers , only

OpenStudy (anonymous):

Do you know how to solve 1.2 * 10 * 10 * 10 ? Let me answer the first = 1200

OpenStudy (phebe):

yess

OpenStudy (anonymous):

Just solve all of those quickly and tell me the answer so I can tell if it is right .

OpenStudy (anonymous):

First solve all of the x 10 ^something , then arrange them . After that rewrite them in their original form

OpenStudy (phebe):

ok

OpenStudy (phebe):

@Imad_Sawal

OpenStudy (anonymous):

yes ?

OpenStudy (anonymous):

I guess there is no need to solve them , you can just arrange the exponents

OpenStudy (phebe):

ok but wat is the answer

OpenStudy (anonymous):

I mean first , 10 ^ 3 , then 10 ^ 3 , then 10 ^ -2 , then 10 ^ -3 .

OpenStudy (anonymous):

You can just type the answer which you think is correct , and I'll ammend and correct it